# Give and Take: An End-To-End Investigation of Giveaway Scam Conversion Rates

### v1

@@ Abstract.
Scams—fraudulent schemes designed to swindle money from victims—have existed for as long as recorded history. However, the Internet’s combination of low communication cost, global reach, and functional anonymity has allowed scam volumes to reach new heights. Designing effective interventions requires first understanding the context: how scammers reach potential victims, the earnings they make, and any potential bottlenecks for durable interventions. In this short paper, we focus on these questions in the context of cryptocurrency giveaway scams, where victims are tricked into irreversibly transferring funds to scammers under the pretense of even greater returns. Combining data from Twitter, [^1] The platform has been known as X since July 2023, but most of our analysis pertains to tweets posted before then so we continue to call it Twitter. YouTube and Twitch livestreams, landing pages, and cryptocurrency blockchains, we measure how giveaway scams operate at scale. We find that 1 in 1000 scam tweets, and 4 in 100,000 livestream views, net a victim, and that scammers managed to extract nearly \$4.62 million from just hundreds of victims during our measurement window.
@@ 1. Introduction
Since the introduction of the Bitcoin whitepaper in 2008, cryptocurrencies and blockchain-based technologies have attracted attention from investors, technologists, researchers, governments, and the general public. Unsurprisingly, this rapid rise in popularity has also attracted scammers seeking to exploit the gap between people’s (high) interest in these technologies and their (low) technical understanding of their operation. Today, there are many types of scams run within the cryptocurrency ecosystem: Ponzi and pyramid schemes (vasek2015freelunch, ; vasek2018ponzi, ; bartoletti2018ponzi, ; chen2018ponzi, ; siu2022bitcointalk, ; kell2023forsage, ); NFT “rug pulls” (roy2023nftscams, ; huang2023rugpulls, ), counterfeiting (das2022nft, ), airdrop scams (roy2023nftscams, ); pump-and-dump schemes (hamrick2018pumpdump, ; xu2019pumpdump, ; xia2020covidscams, ); and giveaway scams (vakilinia2022giveaway, ; li2023giveaway, ; li2023twitter-giveaway, ).
@@ 2. Related Work
Our work builds on a number of prior efforts that have investigated cryptocurrency giveaway scams. Xia et al. (xia2020covidscams, ) document dozens of giveaway scams while searching for Covid-themed cryptocurrency fraud reported by a variety of industry and government sources. Phillips and Wilder (phillips2020tracing, ) analyzed hundreds of giveaways scams (which they refer to as advanced-fee scams) crowdsourced by victims and researchers on CryptoScamDB and URLScan. Li et al. (li2023giveaway, ) deployed the CryptoScamTracker tool, which found thousands of scam domains by examining Certificate Transparency (CT) log data for a six-month period between January and June 2022. The authors used a combination of scam-related keywords to identify potential scam domains, filtering down this list and then validating manually (i.e., that the site in fact hosts a scam) to obtain their scam domain corpus.
@@
In addition to giveaway scams, the security community has explored a range of other cryptocurrency-related scams, including high-yield investment scams (vasek2015freelunch, ; bartoletti2018ponzi, ; vasek2018ponzi, ; boshmaf2020investigating, ; chen2019exploiting, ; jung2019data, ; toyoda2019novel, ; chen2018ponzi, ; siu2023get, ; bartoletti2020dissecting, ; kell2023forsage, ; siu2022bitcointalk, ); fake or high-risk cryptocurrency-related services (e.g., arbitrage bots (li2023towards, ), exchanges (vasek2015freelunch, ; moore2013beware, ; xia2020characterizing, ), and new coins and tokens (gao2020tracking, ; xia2021trade, )); initial coin offering scams (sapkota2020much, ; zetzsche2017ico, ; liebau2019crypto, ; tiwari2020future, ; bian2018icorating, ); sextortion scams (paquet2019spams, ; oggier2020ego, ); technical support scams (acharya2024conning, ); scams and fraud in non-fungible token markets (kshetri2022scams, ; das2022nft, ); and cryptocurrency-themed phishing websites (andryukhin2019phishing, ; holub2018coinhoarder, ; phillips2020tracing, ; he2023txphishscope, ).
@@ 3.1. Data sources
Our Twitter dataset originates from a continuous Internet-wide crawl of public URLs, in a way that respects robots.txt and other rules for crawlers. For the purposes of this study, we restrict our analysis to publicly available tweets. Our dataset covers a period from April 1, 2020 until August 20, 2023. It does not contain every single tweet, and is biased towards tweets that are more discoverable (e.g., posted by accounts with a non-zero number of followers). Access to this data was granted to us only due to the nature of our research project and its potential to disrupt abusive behavior. In total, our snapshot contains tens of billions of tweets posted by hundreds of millions of accounts.
@@
We have access to a proprietary service (chainanalysis, ) that aggregates raw blockchain data and annotates each address indicating (1) the multi-input cluster (reid2013bitcoin, ; meiklejohn2013fistful, ) it belongs to and (2) the likely category of the real-world owner of that cluster. [^2] The service can be configured to provide the name of the cluster owner, but our agreement meant we had access only to the broader category. These latter annotations are obtained via regular transactions that this company conducts with known cryptocurrency services, such as exchanges, and the former annotation avoids false positives (such as those induced by Coinjoin transactions) using proprietary heuristics.
@@ 3.2. Identifying giveaway livestreams
The measurement pipeline frequently polled available livestreams via platform APIs, recorded them, crawled the pages promoted by the QR codes and links, and annotated each scam page domain name with blockchain metadata. The design decisions behind this system are based on a 14-day pilot study conducted July 1–14, 2023. While we crawled both YouTube and Twitch, we did not detect any giveaway scams during our collection window for Twitch despite past reports of such scams (twitch-scams2022, ). As a result, we focus our subsequent discussion and analysis on YouTube, deferring Twitch-specific details to Appendix B.1.
@@
The measurement pipeline starts data collection by periodically querying the YouTube API for available streams and associated metadata. Given a set of keywords, the API returns livestreams that are associated with the keywords (in a manner similar to Google Search). Our system queried the API at the maximum rate allowed, which meant retrieving the list of livestreams and channel metadata (e.g., number of subscribers) every 30 minutes and retrieving livestream metadata (e.g., concurrent and total viewers) every 7.5 minutes. We derived our set of search keywords using a combination of popular cryptocurrency names and those used by Li et al. (li2023giveaway, ) (more details in Appendix B.2).
@@
Stream and chat recording
The system recorded each livestream (using Streamlink (Streamlink:online, )) and its associated chat (using the YouTube API) to identify linked URLs. In our pilot study, we observed that URLs either appeared in chat or were embedded in the video using a QR code. To capture embedded QR codes, the system recorded videos for two seconds at a time, with snapshots taken every 7.5 minutes. For chats, the system polled the chat API every 7.5 minutes to obtain the last 70 historical messages (the maximum set by YouTube). Appendix B.3 provides more information on why two-second samples and 70 historical messages are sufficient.
@@
The system crawled each potential scam URL, revisiting both old and newly discovered websites daily until (1) we reached the end of our collection period, or (2) fetching the URL resulted in an error three days in a row. Prior work (li2023giveaway, ) suggested that some giveaway scam sites adopted anti-bot tactics. We identified four types of cloaking behavior in our pilot study (Appendix B.4) and counter these behaviors by: (1) using a VPN to mask the system IP address; (2) spoofing the User-Agent string to appear as a popular device and browser; (3) having a heuristic module to click through common front pages; and (4) verifying the crawler with Cloudflare to bypass anti-bot detection.
@@
We ran our measurement pipeline from July 24, 2023 to January 21, 2024 (26 weeks), identifying a total of 2,069 livestreams from 1,632 different channels (Table 1). These streams linked to 343 distinct domains. Our data is continuous throughout this period apart from 11 days where we experienced infrastructure outages. [^4] These dates are: 2023-08-15, 2023-08-16, 2023-09-01, 2023-09-28, 2023-10-06, 2023-11-18, 2023-11-19, 2023-12-26, 2023-12-12, 2024-01-06, and 2024-01-21.
@@ 3.3. Limitations
As with any measurement study, there are a number of limitations with our methodology. Our Twitter dataset may omit some tweets, resulting in us underestimating the volume of scam tweets or the revenue generated by scams; likewise, this may happen due to missing or inaccurate cryptocurrency addresses in the CryptoScamTracker dataset (as indeed we observed manually). Given that only 9% of the domains previously identified by CryptoScamTracker overlap with Twitter, it is also clear that many scams are promoted via other distribution channels—be it social media, email, or messaging. Finally, our keyword-based matching strategies for YouTube and Twitch likely resulted in omitting some scams. Despite these limitations, we are able to draw insights into the promotion of giveaway scams and provide the first estimates on the conversion rate and revenue of such campaigns.
@@ 4.1. Volume
In total, we observed 457,248 scam tweets and 2,069 scam livestreams. We present a weekly timeline of scam activity for Twitter in Figure 3 and YouTube in Figure 4. While scam lures appeared consistently throughout our measurement window, there are notable bursts of activity. For Twitter, we saw a single peak in activity in March 2022, with scammers posting a maximum of 90,984 tweets in a single week. On YouTube, we observed a burst in September 2023 and again over the holiday period of December–January 2024, peaking at 289 streams and 1,869,399 views in a single week. While these scams are smaller in scale compared to spam email campaigns involving tens of millions of messages (kanich2008spamalytics, ), as we show shortly, they are nevertheless highly profitable.
@@ 4.2. Discoverability
To reach a broad audience on Twitter, scammers hashtagged their tweets 96% of the time (e.g., using a cryptocurrency’s name or ticker symbol). This meant that users searching for the latest news related to popular coins (e.g., #btc, #eth) might then interact with the scammer’s tweets. While scammers also pursued alternative audience-generating tactics on Twitter, such as mentioning users or replying to popular tweets, these strategies occurred in only 0.1% and 0.3% of tweets respectively.555 For Twitter, our dataset lacks any information on the number of views per tweet, or the following or follower counts of accounts. As such, we focus on discoverability. For YouTube, scammers reached victims using a combination of organic viewership and search discoverability. Channels used by scammers to host livestreams had a median of 16.8K subscribers, with the single largest channel having 19 million subscribers—liked a compromised pre-existing channel. Beyond existing audiences, 93% of livestreams included a cryptocurrency keyword in their title, their description, or the name of the associated channel.
@@ 5.1. Isolating relevant payments
We focused on the three most popular coins targeted by scammers (per Section 4.3): Bitcoin (BTC), Ethereum (ETH), and Ripple (XRP). For each coin, we identified all incoming transactions to cryptocurrency addresses that appeared on a giveaway scam landing page posted to Twitter or YouTube. We filtered the set of incoming transactions to an identified scam address in two key ways. First, following Gomez et al. (gomez2023cybercrime, ), we ignored transactions where the sender is a known scam address, as this clearly represents the consolidation of funds rather than a payment from an actual victim. Second, we focused on payments that co-occurred with a lure in order to capture the lure’s effectiveness in driving victims to the scam. We acknowledge that the resulting set of payments might both include non-victim transactions (from unknown scammer addresses) and exclude victim transactions (due to lures that we were not aware of); nevertheless, we use it as an approximation of victim payments to the scam. We believe this is a more accurate estimate than previous studies, which considered all incoming transactions (li2023giveaway, ), but also compare with the total revenue (across all transactions) and leave as future work a further refinement of our analysis that takes other behaviors into account.

### v2

@@ Abstract.
The Internet’s combination of low communication cost, global reach, and functional anonymity has allowed fraudulent scam volumes to reach new heights. Designing effective interventions requires first understanding the context: how scammers reach potential victims, the earnings they make, and any potential bottlenecks for durable interventions. In this short paper, we focus on these questions in the context of cryptocurrency giveaway scams, where victims are tricked into irreversibly transferring funds to scammers under the pretense of even greater returns. Combining data from Twitter, [^1] The platform has been known as X since July 2023, but most of our analysis pertains to tweets posted before then so we continue to call it Twitter. YouTube and Twitch livestreams, landing pages, and cryptocurrency blockchains, we measure how giveaway scams operate at scale. We find that 1 in 1000 scam tweets, and 4 in 100,000 livestream views, net a victim, and that scammers managed to extract nearly \$4.62 million from just hundreds of victims during our measurement window.
Giveaway Scams; Blockchain Security; Internet Security
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM Internet Measurement Conference; November 4–6, 2024; Madrid, Spain.booktitle: Proceedings of the 2024 ACM Internet Measurement Conference (IMC ’24), November 4–6, 2024, Madrid, Spainisbn: 979-8-4007-0592-2/24/11doi: 10.1145/3646547.3689005ccs: Security and privacy Social network security and privacy
@@ 1. Introduction
Since the introduction of the Bitcoin whitepaper in 2008, cryptocurrencies and blockchain-based technologies have attracted attention from investors, technologists, researchers, governments, and the general public. Unsurprisingly, this rapid rise in popularity has also attracted scammers seeking to exploit the gap between people’s (high) interest in these technologies and their (low) technical understanding of their operation. Today, there are many types of scams run within the cryptocurrency ecosystem: Ponzi and pyramid schemes (vasek2015freelunch, ; vasek2018ponzi, ; bartoletti2018ponzi, ; chen2018ponzi, ; siu2022bitcointalk, ; kell2023forsage, ); NFT “rug pulls” (roy2023nftscams, ; huang2023rugpulls, ), counterfeiting (das2022nft, ), airdrop scams (roy2023nftscams, ); pump-and-dump schemes (dhawan2023pumpdump, ; hamrick2018pumpdump, ; kamps2018pumpdump, ; xia2020covidscams, ; xu2019pumpdump, ); and giveaway scams (vakilinia2022giveaway, ; li2023giveaway, ; li2023twitter-giveaway, ).
@@ 2. Related Work
Our work builds on a number of prior efforts that have investigated cryptocurrency giveaway scams. Xia et al. (xia2020covidscams, ) document dozens of giveaway scams while searching for Covid-themed cryptocurrency fraud reported by a variety of industry and government sources. Phillips and Wilder (phillips2020tracing, ) analyzed hundreds of giveaways scams (which they refer to as advanced-fee scams) crowdsourced by victims and researchers on CryptoScamDB and URLScan. Li et al. (li2023giveaway, ) deployed the CryptoScamTracker tool, which found thousands of scam domains by examining Certificate Transparency (CT) data for a six-month period between January and June 2022. The authors used a combination of scam-related keywords to identify likely scam domains, filtering down this list and then validating manually (i.e., that the site does host a scam) to obtain their scam domains.
@@
In addition to giveaway scams, the security community has explored a range of other cryptocurrency-related scams, including high-yield investment scams (vasek2015freelunch, ; bartoletti2018ponzi, ; vasek2018ponzi, ; boshmaf2020investigating, ; chen2019exploiting, ; jung2019data, ; toyoda2019novel, ; chen2018ponzi, ; siu2023get, ; bartoletti2020dissecting, ; kell2023forsage, ; siu2022bitcointalk, ); fake or high-risk cryptocurrency-related services (e.g., arbitrage bots (li2023towards, ), exchanges (vasek2015freelunch, ; moore2013beware, ; xia2020characterizing, ), pump-and-dump scams (dhawan2023pumpdump, ; hamrick2018pumpdump, ; kamps2018pumpdump, ; xia2020covidscams, ; xu2019pumpdump, ); and new coins and tokens (gao2020tracking, ; xia2021trade, )); initial coin offering scams (sapkota2020much, ; zetzsche2017ico, ; liebau2019crypto, ; tiwari2020future, ; bian2018icorating, ); sextortion scams (paquet2019spams, ; oggier2020ego, ); technical support scams (acharya2024conning, ); scams and fraud in non-fungible token markets (kshetri2022scams, ; das2022nft, ); and cryptocurrency-themed phishing websites (andryukhin2019phishing, ; holub2018coinhoarder, ; phillips2020tracing, ; he2023txphishscope, ). In contrast to scams that require more sophisticated methods for turning a profit — price manipulation (e.g., pump-and-dump), hierarchies of participants to asymmetrically channel funds (e.g., Ponzi schemes), etc. — giveaway scams are quite straightforward. Victims simply give their funds to scammers.
@@ 3.1. Data sources
Our Twitter dataset originates from Google’s Internet-wide crawl of public URLs, which respects robots.txt and other rules for crawlers. This means our analysis considers only publicly available tweets. Our dataset covers a period from April 1, 2020 until August 20, 2023. It does not contain every single tweet, and is biased towards tweets that are more discoverable (e.g., posted by accounts with a non-zero number of followers). Within Google, access to this data was granted to us only due to the nature of our research project and its potential to disrupt abusive behavior. In total, our snapshot contains tens of billions of tweets posted by hundreds of millions of accounts.
@@
We have access to the Chainalysis Crypto Investigations tool (chainanalysis, ), which processes raw blockchain data and annotates each address indicating (1) the multi-input cluster (reid2013bitcoin, ; meiklejohn2013fistful, ) it belongs to, meaning the cluster of addresses that certain heuristics indicate are all operated by the same entity, and (2) the likely category (e.g., exchange, mixer, token smart contract) of the real-world operator of that cluster. [^2] The tool can be configured to provide the name of the cluster operator, but our agreement with Chainalysis meant we had access only to the broader category. These latter annotations are obtained via regular transactions that Chainalysis conducts with known cryptocurrency services, such as exchanges, and the former annotation avoids false positives (such as those induced by Coinjoin transactions (coinjoin, )) using proprietary heuristics. While there is some potential for error in both of these areas, Chainalysis data has been used in previous academic studies (portnoff2017backpage, ; huang2018tracking, ; kappos2022peel, ) and was recently ruled to be sufficiently reliable as to be admissible evidence in court (chainalysis-daubert, ).
@@ 3.2. Identifying giveaway livestreams
Our pipeline polled available livestreams via platform APIs, recorded them, crawled the pages promoted by the QR codes and links, and annotated each scam domain with blockchain metadata. The design decisions are based on a 14-day pilot study conducted July 1–14, 2023. While we crawled both YouTube and Twitch, we did not detect any giveaway scams during our collection window for Twitch despite past reports of such scams (twitch-scams2022, ). As a result, we focus our subsequent discussion and analysis on YouTube, deferring Twitch-specific details to Appendix B.1.
@@
Our pipeline starts data collection by periodically querying the YouTube API for available streams and associated metadata. Given a set of keywords, the API returns livestreams that are associated with the keywords (in a manner similar to Google Search). Our system queried the API at the maximum rate allowed, which meant retrieving the list of livestreams and channel metadata (e.g., number of subscribers) every 30 minutes and retrieving livestream metadata (e.g., concurrent and total viewers) every 7.5 minutes.
We derived our set of search keywords using a combination of popular cryptocurrency names and those used by Li et al. (li2023giveaway, ). In more detail, we opted for a set of keywords consisting of the names and ticker symbols of the top 20 cryptocurrencies (as listed on coinmarketcap.com on July 1, 2023), [^3] In three cases (for ADA, SOL, and DOT), we appended the word “coin” to the ticker symbol to avoid introducing irrelevant streams. along with the keywords used by CryptoScamTracker (li2023giveaway, , Table XI) and two catch-all terms: “cryptocurrency” and “crypto.” We discuss the effectiveness of these keywords in more detail in Appendix B.2.
Stream and chat recording
The system recorded each livestream (using Streamlink (Streamlink:online, )) and its associated chat (using the YouTube API) to identify linked URLs. Our final pipeline recorded videos for two seconds at a time, with snapshots taken every 7.5 minutes, to capture QR codes embedded in the video. For chats, it polled the chat API every 7.5 minutes to obtain the last 70 historical messages (the maximum set by YouTube).
We chose these parameters because in our pilot study we observed that URLs either appeared in chat or were embedded in the video using a QR code. We thus considered (1) how long a QR code lasts once it appears and (2) how many chat messages are posted during the lifespan of a scam livestream. We started by running our pipeline at maximum capacity on YouTube, with the default setup of recording a stream for two seconds as well as retrieving up to 70 historical chat messages every 7.5 minutes. In addition, for a subset of the livestreams, we kept recording the stream upon detecting a QR code. With this setup, we manually identified 276 streams that promoted 59 unique giveaway websites. We first examined how long a QR code lasted after we first detected it. For 41 scam livestreams, we continued to record them after detecting a QR code. We found that for a majority of the livestreams, a QR code remained displayed continuously for a reasonable amount of time (mean of 7,200 seconds and median of 3,140 seconds), except for one case where the QR code was displayed periodically for around 15 seconds. This finding suggests that two-second samples are enough to capture a QR code in the livestream, if present. Second, we examined scam streams that had chat messages. We found that all scam streams had few chat messages (less than 10) and no user interactions, and thus decided that 70 historical chat messages every 7.5 minutes was sufficient.
@@
The system crawled each potential scam URL, revisiting both old and newly discovered websites daily until (1) we reached the end of our collection period, or (2) fetching the URL resulted in an error three days in a row. Prior work (li2023giveaway, ) suggested that some giveaway scam sites adopted anti-bot tactics. We identified four types of cloaking behavior in our pilot study: (1) IP-based cloaking, in which requests from an institutional network resulted in 403 responses but ones from a residential network did not; (2) user agent-based cloaking, in which requests from browsers that were not running on Windows or Mac resulted in 403 responses; (3) customized front pages, in which human interaction was required to click on a button or select a desired cryptocurrency; and (4) anti-bot detection through Cloudflare. We counter these behaviors by: (1) using a VPN to mask the system IP address; (2) spoofing the User-Agent string to appear as a popular device and browser; (3) having a heuristic module to click through common front pages; and (4) verifying the crawler with Cloudflare to bypass anti-bot detection. [^4] https://blog.cloudflare.com/friendly-bots
@@
We ran our measurement pipeline from July 24, 2023 to January 21, 2024 (26 weeks), identifying a total of 2,069 livestreams from 1,632 different channels (Table 1). These streams linked to 343 distinct domains (all but one domains met criteria 2 and 303 domains met criteria 3). Our data is continuous throughout this period apart from 11 days of infrastructure outages. [^6] These dates are: 2023-08-15, 2023-08-16, 2023-09-01, 2023-09-28, 2023-10-06, 2023-11-18, 2023-11-19, 2023-12-26, 2023-12-12, 2024-01-06, and 2024-01-21.
@@ 3.3. Limitations
As with any measurement study, there are a number of limitations with our methodology. Our Twitter dataset may omit some tweets, resulting in us underestimating the volume of scam tweets or the revenue generated by scams; likewise, this may happen due to missing or inaccurate cryptocurrency addresses in the CryptoScamTracker dataset (as indeed we observed manually). Given that only 9% of the domains previously identified by CryptoScamTracker overlap with Twitter, it is also clear that many scams are promoted via other distribution channels—be it social media, email, or messaging. Similarly, our keyword-based matching strategies for YouTube and Twitch likely resulted in omitting some scams. Finally, our Twitter and YouTube data are drawn from non-overlapping periods of time (early 2022 vs. late 2023), which means the tactics of scammers may have evolved so we cannot directly compare profitability. Despite these limitations, we are able to draw insights into the promotion of giveaway scams and provide the first estimates on the conversion rate and revenue of such campaigns.
@@ 4.1. Volume
In total, we observed 457,248 scam tweets and 2,069 scam livestreams. We present a weekly timeline of scam activity for Twitter in Figure 3 and YouTube in Figure 4. While scam lures appeared consistently throughout our measurement window, there are notable bursts of activity. Twitter had a single peak in activity in March 2022, with scammers posting a maximum of 90,984 tweets in a single week. YouTube had a burst in September 2023 and again over the holiday period of Dec–Jan 2024, peaking at 289 streams and 1,869,399 views in a single week. While these scams are smaller in scale compared to spam email campaigns involving tens of millions of messages (kanich2008spamalytics, ), as we show shortly, they are nevertheless highly profitable.
@@ 4.2. Discoverability
To reach a broad audience on Twitter, scammers hashtagged their tweets 96% of the time (e.g., using a cryptocurrency’s name or ticker symbol). Users searching for the latest news related to popular coins (e.g., #btc, #eth) might then interact with the scammer’s tweets. While scammers also pursued alternative tactics on Twitter, such as mentioning users or replying to popular tweets, these strategies occurred in only 0.1% and 0.3% of tweets respectively.777 For Twitter, our dataset lacks any information on the number of views per tweet, or the following or follower counts of accounts. As such, we focus on discoverability. For YouTube, scammers reached victims using a combination of organic viewership and search discoverability. Channels used by scammers to host livestreams had a median of 16.8K subscribers, with the single largest channel having 19 million subscribers—likely a compromised pre-existing channel. Beyond existing audiences, 93% of livestreams included a cryptocurrency keyword in their title, their description, or the name of the associated channel.
@@ 5.1. Isolating relevant payments
We focused on the three most popular coins targeted by scammers (per Section 4.3): Bitcoin (BTC), Ethereum (ETH), and Ripple (XRP). For each coin, we identified all incoming transactions to cryptocurrency addresses that appeared on a giveaway scam landing page posted to Twitter or YouTube. We filtered the set of incoming transactions to an identified scam address in two key ways. First, following Gomez et al. (gomez2023cybercrime, ), we ignored transactions where the sender is a known scam address since this represents the consolidation of funds rather than a payment from an actual victim. Second, we focused on payments that co-occurred with a lure to capture the lure’s effectiveness in driving victims to the scam. We acknowledge that the resulting set of payments might both include non-victim transactions (from unknown scammer addresses) and exclude victim transactions (due to lures not in our data); nevertheless, we use it as an approximation of victim payments to the scam. We believe this estimate is more accurate than previous studies, which considered all incoming transactions (li2023giveaway, ), but also compare with the total revenue (across all transactions) and leave as future work refinements to our analysis that take other behaviors into account.